# Sandwich GAN: Image Reconstruction from Phase Mask based Anti-dazzle Imaging

### 24 Feb 2024 v1

@@ Abstract
Conventional camera systems are susceptible to the adverse effects of laser dazzle, which may over-saturate an image or cause permanent damage to pixels. To address this problem, we developed an approach combining point spread function engineering whereby a wavefront-coded mask in the pupil plane blurs both the laser and scene, together with a deep neural sandwich network. In addition to protecting the sensor, our approach jointly removes the laser from the scene and reconstructs a satisfactory deblurred image. Image recovery is achieved by wrapping two generative adversarial networks (GANs) around a learnable non-blind image deconvolution module. We trained the Sandwich GAN (SGAN) to suppress the peak laser irradiance as high as $10^{6}$ times the sensor saturation threshold - the point at which the bare system without the phase mask may exhibit damage. The end-to-end training includes physics-based modeling of the imaging system whereby a laser having an arbitrary angle of incidence is superimposed on images from a large publicly available library. The trained system was validated in the laboratory for laser strengths up to $10^{4}$ times the saturation value. The proposed image restoration model quantitatively and qualitatively outperforms other methods for a wide range of scene contents, illumination conditions, laser strengths, and noise characteristics.
@@
Computational imaging, Diffractive mask, Image reconstruction, Deep Learning, Generative Adversarial Networks
@@ 1 Introduction
Our approach, illustrated in Figure 1 is an example of computational imaging whereby a diffractive phase mask is introduced into the optical path (Figure 1(a)) so that the ground truth scene is encoded in the recorded image in such a way to facilitate a reconstructed image. The phase mask is engineered to diffuse the laser light at the pupil plane, thereby reducing the peak irradiance of the focal spot while simultaneously providing adequate contrast across the transmitted spatial frequencies of the system [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. The first task protects the sensor from laser damage and the second ensures satisfactory image quality. The laser, represented by a uniform planar wave front of wavelength $\lambda$, is assumed to overfill the aperture of the imaging lens at an arbitrary angle of incidence. Here we assume that the background scene is incoherently illuminated with quasi-monochromatic light centered at the same wavelength. The phase mask, placed adjacent to the imaging lens, transforms the uncoded point spread function (PSF) into a coded PSF. Whereas the uncoded PSF produces high irradiance across a few pixels, the coded PSF spreads the laser power over many hundreds of pixels, resulting in a significantly smaller peak irradiance value. The insertion of the coded phase mask blurs the scene and introduces a laser diffraction pattern on the recorded image, depicted in Figure 1(b).
@@
Figure 1: Anti-dazzle imaging at a glance. (a) Schematic of the wavefront coded imaging system to protect a sensor from laser irradiation. Quasi-monochromatic background illumination $\lambda_{b}$, which is close to the laser wavelength $\lambda_{b}$ is assumed. A five half-ring phase mask [38] is placed adjacent to the lens. (b) Simulation of a phase coded image of far-field background scene and potential sensor damaging laser. (c) Restored image using our SGAN-F model and (d) Ground truth image.
@@
While the coded phase mask protects the sensor from being damaged by laser radiation, we face several challenges in achieving image restoration. Compared to an unprotected system, coded PSF may lead to significant image blur and saturation, as well as the lose of image boundary information in recorded images. Although the PSF may be determined via simulation or measurement, the use of a single image restoration algorithm may not be sufficient to produce desirable results due to combined degradations as well as unknowns such as illumination conditions and the variable nature of the laser strength and direction. With this work, we introduce a Sandwich Generative Adversarial Networks (SGAN), which wraps a learnable nonblind deconvolution module between two GANs, for the image restoration a phase coded anti-dazzle imaging system. The model inpaints the laser-induced saturation, outpaints the outer lost boundary area, and reduces sensor noise through a self-attention conditional GAN, thereby producing a pre-restored image. To remove image blur caused by the use of phase mask, a learnable deconvolution is applied to a set of extracted features of the pre-restored result. The deblurred features are combined and refined with a multiscale self-attention GAN to generate the final restored image. Three variants of the SGAN model are investigated. They include the basic model SGAN-B, the enhanced model SGAN-E, and the frequency model SGAN-F. These three models make use of different loss functions and feature representations in neural networks. High-fidelity images may be reconstructed using the SGAN-F model (see Figure 1(c)), which resembles well the ground truth shown in Figure 1(d). The specific contributions of this work are summarized as follows.
@@
  • A sandwich neural network architecture that combines discriminative and generative image reconstruction methods is tailored to a wavefront coded anti-dazzle imaging system. End-to-end training encourages global optimization of the model, while the utilization of high-performing losses and Fourier features in the neural networks further improves the image restoration accuracies. The state-of-the-art performance is validated on both numerical and experimental data.
@@
  • Physics-based modeling of the imaging pipeline is formulated, which characterizes the experimental system. A large set of training data is generated numerically. Compared to experimental acquisitions, numerical modeling obviates both hardware calibrations (e.g., image registration) and data augmentation (e.g., laser parameters, ambient illumination, and noise). Ground truth images that formed at intermediate stages that otherwise may not be able to measure in practice (e.g., the uncropped irradiance maps) are also numerically generated to supervise the neural network training.
@@ 2 Related Work
Coded Aperture. Computational imaging and deep optics [45, 46] is an emerging area that focuses on improving and extending the capability of traditional imaging and camera systems using optical and computational methods. By altering light transmission at the pupil plane using an amplitude mask or a phase mask, the coded aperture approach has been investigated in many applications, such as achromatic imaging[47, 48], motion deblur [49], extended depth of field [50, 51, 52], lens glare suppression [53, 54, 55], high dynamic range imaging [56, 57], light field imaging [58], privacy-preserving imaging [59, 60], hyperspectral imaging [61, 62], interferemtric imaging [63, 64, 65], compressive sensing [66, 67], holographic imaging [68, 69], super-resolved imaging [70, 71], and lensless imaging [72]. An amplitude mask modulates the incident light by partially blocking it with a binary pattern, and the PSF is the shadow of the mask. The wavefront-coded aperture bends the incident light using a transparent phase mask, and the PSF is governed by its height profile. Compared to amplitude masks, phase elements allow higher light throughput and signal-to-noise ratio, and finer light modulation.
@@
where $h$ is the PSF of a shift-invariant system and the operator *** represents the spatial convolution. For differentiable regularization terms, the solution to Eq. 2 may be estimated by analytical inversions, such as inverse filters, Wiener deconvolution for Tikhonov regularization [73], and fast deconvolution for hyper-Laplacian prior [74]. Straightforward deconvolution methods were used in early coded aperture systems [49, 53, 58, 75]. These approaches are sensitive to noise, and restored images may suffer from ringing artifacts. Alternatively, regularization terms are modeled as a known statistical distribution (e.g., heavy-tailed distribution [76], hyper-Laplacian [77], sparse gradients [78], etc.) or are learned through a shrinkage function [79, 80]. Iterative solutions may be derived using splitting methods, such as half-quadratic splitting [81], reweighted least square [82], primal-dual[83], and alternating direction multiplier method [84]. Iterative methods have been applied in many coded aperture systems [85, 86, 87, 88].
@@
Learning a direct mapping from $s$ to $\hat{b}$ using DNN has been extensively studied for many low-level computer vision tasks [105], including image deblurring, denoising, deraining, dehazing, in- and outpainting, as well as image restoration from low light conditions. Most vision models have an encoder-decoder framework, including UNet[106], residual network [107], and recurrent network [108]. The self-attention mechanism [109] seeks to capture long-range image dependencies. It serves as the building blocks of vision transformers [110, 111, 112]. Its quadratic computational complexities can be reduced to linear using window partitioning techniques [113, 114] or global approximation [115]. Self-attention has also been approximated by the spectral power density through the efficient fast Fourier transform (FFT) [116]. As an alternative, the spatial gating unit [117] was introduced with the MLP to achieve a performance comparable to that of transformers, where similar block partition and axis swap techniques were used to reduce its complexity [118]. Additionally, changing the image basis through linear or nonlinear transforms to better suit the orthogonally preconditioned optimizers (e.g., ADAM, ADAgrad, etc.) was found to accelerate the convergence and improve the performance of the DNNs. One notable finding involves the use of FFT features to correct spectral bias and improve the learning of high-frequency functions from low dimension data [119, 120, 121]. Another important achievement is the use of feature matching (or perceptual) loss [122] in the image intensity space, where features are extracted by pretrained VGG Net [123] or discriminators [124]. These feature representations and losses have been shown to increase the accuracy of image restoration [125, 126, 118, 127]. Furthermore, by breaking down a complicated task into subproblems and solving them progressively, multistage and multiscale frameworks allow supervision and feature fusion in multiple restoration stages and multiple image scales[128, 129] and encourage the recovery of image details. Embedding of kernel functions [130] and image coordinates [131, 132, 133] into neural networks has also been introduced, respectively, to utilize the knowledge of system PSF and positional information of an image.
@@ 4.1 Pre-Restoration with Conditional GAN
Here a residual UNet is designed for $G_{1}$, and the Markovian discriminator (PatchGAN) [146] is employed for $D_{1}$ in our SGAN-B/E models. The double convolutional layers of the original UNet are replaced with residual blocks [107]. The neural net blocks of this module are listed in Table II. For the SGAN-F model, we make use of the residual FFT (RFT) blocks [127] in $G_{1}$ and FFT convolutional (FTC) blocks [120] in $D_{1}$ respectively. The structure of the net blocks and the values of the hyperparameters $\lambda_{ADV}$ and $\lambda_{GP}$ are provided in the Supplementary Information.
@@ 6.1 Baseline Methods and Evaluation Metrics
For the anti-dazzle imaging, FFT representations and losses are found to be highly effective in restoring high-frequency image details and producing sharper restored images. SSIM [173] and multiscale discriminator feature matching (MDF) losses [124] appear to have similar but weaker impacts on image structures restoration. L1 loss is shown to encourage in- and outpainting of missing areas, and VGG loss [122] improves perceptual qualities. Here we categorize the models as low- and high-performing based on the loss functions. The low-performing models, which trained with combinations of L1, SSIM, and MDF losses, involve MPRNet [129], Pix2Pix [146], ST-CGAN [172], Uformer [111], DeepWiener [92], WienerNet [91], and our SGAN-B. The high-performing models, which make use of FFT and VGG losses, include DeblurGAN [152], Stripformer [126], MIMO-UNet [125], MAXIM-2S [118], DeepRFT [127], and our SGAN-E/F. The use of FFT features in DeepRFT [127] and our SGAN-F further improves their performance in image restoration.
@@ 6.6 Discussions
Future work is required to extend this approach to broadband illumination and a variety of laser wavelengths. Examination of other unfavorable imaging conditions, such as shift-variant distortions, motion blur, atmosphere, and adverse weathers, are also indispensable towards practical applications. Taking advantage of spectral representations in neural networks may be worth further investigation to develop fully blind image restoration without compromising the quality of the restored images. Even higher-fidelity reconstruction may be achieved by mitigating the aliasing issue presented in multiscale training using learnable subpixel sampling techniques [179]. Provided limited computational resources, deep learning based super-resolution may allow restoration of high-resolution images from degraded inputs of low resolutions. Furthermore, a fully differentiable end-to-end imaging pipeline [180] that jointly optimizes the phase function and image restoration algorithm may further improve the performance of the antidazzle imaging system.

### 04 Mar 2024 v2

# Learning to See Through Dazzle

@@ Abstract
Machine vision is susceptible to laser dazzle, where intense laser light can blind and distort its perception of the environment through oversaturation or permanent damage to sensor pixels. Here we employ a wavefront-coded phase mask to diffuse the energy of laser light and introduce a sandwich generative adversarial network (SGAN) to restore images from complex image degradations, such as varying laser-induced image saturation, mask-induced image blurring, unknown lighting conditions, and various noise corruptions. The SGAN architecture combines discriminative and generative methods by wrapping two GANs around a learnable image deconvolution module. In addition, we make use of Fourier feature representations to reduce the spectral bias of neural networks and improve its learning of high-frequency image details. End-to-end training includes the realistic physics-based synthesis of a large set of training data from publicly available images. We trained the SGAN to suppress the peak laser irradiance as high as $10^{6}$ times the sensor saturation threshold - the point at which camera sensors may experience damage without the mask. The trained model was evaluated on both a synthetic data set and data collected from the laboratory. The proposed image restoration model quantitatively and qualitatively outperforms state-of-the-art methods for a wide range of scene contents, laser powers, incident laser angles, ambient illumination strengths, and noise characteristics.
@@
Laser dazzle, Phase mask, Physics-based image synthesis, Computer vision, Low-level vision, Image restoration, Image inpainting, Image deblurring, Image denoising, Low-light image enhancement, Computational imaging, Machine learning
@@ 1 Introduction
Our approach, illustrated in Figure 1 is an example of computational imaging whereby a diffractive phase mask is introduced into the optical path (Figure 1(a)) so that the ground truth scene is encoded in the recorded image in such a way to facilitate a restored image. The first task protects the sensor from laser damage and the second ensures satisfactory image quality. The phase mask is engineered to diffuse the laser light at the pupil plane, thereby reducing the peak irradiance of the focal spot while simultaneously providing adequate contrast across the transmitted spatial frequencies of the system [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. The laser, represented by a uniform planar wave front of wavelength $\lambda$, is assumed to overfill the aperture of the imaging lens at an arbitrary angle of incidence. Here we consider that the background scene is incoherently illuminated with quasi-monochromatic light centered at the same wavelength. The phase mask, placed adjacent to the imaging lens, transforms the uncoded point spread function (PSF) into a phase-coded PSF. Whereas the uncoded PSF produces high irradiance across a few pixels, the phase-coded PSF spreads the laser power over many hundreds of pixels, resulting in a significantly lower peak irradiance value. The insertion of the coded phase mask blurs the scene and introduces a laser diffraction pattern on the recorded image, depicted in Figure 1(b).
@@
Figure 1: Laser dazzle protection at a glance. (a) Schematic of the wavefront-coded camera to diffuse laser irradiation. Quasi-monochromatic background illumination $\lambda_{b}$, which is close to the laser wavelength $\lambda_{b}$ is assumed. A five half-ring phase mask [38] is placed adjacent to the lens. (b) Simulation of a phase-coded image of a background scene and potential sensor-damaging laser. (c) Restored image using our SGAN-F model and (d) Ground truth image.
@@
While the coded phase mask protects the sensor from being damaged by laser radiation, we face several challenges in achieving high-quality image restoration. Compared to an unprotected system, coded PSF may lead to significant image blur and saturation, as well as loss of image boundary information in recorded images. Although the PSF may be determined via simulation or measurement, the use of a single image restoration algorithm may not be sufficient to produce desirable results due to complex degradations as well as unknowns such as illumination conditions and the variable nature of the laser strength and direction. With this work, we introduce a Sandwich Generative Adversarial Network (SGAN), which wraps a learnable non-blind deconvolution module between two GANs, for the image restoration. The neural network model inpaints the laser-induced saturation, outpaints the lost image boundary areas, and reduces sensor noise through a self-attention conditional GAN, thereby producing a pre-restored image. To remove image blur caused by the use of phase mask, a learnable deconvolution is applied to a set of extracted features of the pre-restored result. The deblurred features are combined and refined with a multiscale self-attention GAN to generate the final restored image. Three variants of the SGAN model are investigated. They include the basic model SGAN-B, the enhanced model SGAN-E, and the frequency model SGAN-F. These three models make use of different feature representations and loss functions in their respective neural networks. A high-fidelity image restored by our SGAN-F is presented in Figure 1(c), which resembles the ground truth image shown in Figure 1(d). The specific contributions of this work are summarized as follows.
@@
  • A sandwich neural network architecture is introduced to restore images from laser-dazzled images, where a phase mask is employed to protect the camera from being damaged. The model combines discriminative and generative methods to solve varying image degradations. End-to-end training encourages global optimization of the model, while the utilization of Fourier features and high-performing losses in the neural network reduces its spectral bias and further improves the image restoration accuracies. The state-of-the-art performance is validated on both numerically synthesized and experimentally acquired data.
@@
  • Physics-based image synthesis is formulated, characterizing the real-life imaging pipeline. A large set of training images is generated numerically. Compared to experimental acquisitions, numerical synthesis obviates both hardware calibrations (e.g., image registration) and data augmentation (e.g., laser parameters, ambient illumination, and noise). Ground truth images that formed at intermediate stages and otherwise may not be able to measure in practice (e.g., the uncropped irradiance maps) are also numerically generated to supervise the neural network training.
@@ 2 Related Work
Coded Aperture. Computational imaging and photography are emerging areas that focus on improving and extending the capabilities of traditional imaging and camera systems using optical and computational methods [44, 45]. By altering light transmission at the pupil plane using an amplitude mask or a phase mask, the coded aperture approach has been investigated in many applications, such as coded exposure [46], achromatic imaging[47, 48], high dynamic range [49, 50], lens glare suppression [51, 52], extended depth of field [53, 54, 55], light field imaging [56], lensless imaging [57], privacy-preserving imaging [58, 59], hyperspectral imaging [60, 61], holography [62, 63], compressive sensing [64, 65], super-resolved imaging [66, 67], and interferemtric imaging [68, 69, 70]. An amplitude mask modulates the incident light by partially blocking it with a binary pattern, and the PSF is the shadow of the mask. The wavefront-coded aperture bends the incident light using a transparent phase mask, and the PSF is governed by its height profile. Compared to amplitude masks, phase elements allow higher light throughput and signal-to-noise ratio, and finer light modulation.
@@
where $h$ is the PSF of a shift-invariant system and the operator *** represents the spatial convolution. For differentiable regularization terms, the solution to Eq. 2 may be estimated by analytical inversions, such as inverse filters, Wiener deconvolution for Tikhonov regularization [71], and fast deconvolution for hyper-Laplacian prior [72]. Straightforward deconvolution methods were used in early coded aperture systems [46, 73, 56, 74]. These approaches are sensitive to noise, and restored images may suffer from ringing artifacts. Alternatively, regularization terms are modeled as a known statistical distribution (e.g., heavy-tailed distribution [75], hyper-Laplacian [76], sparse gradients [77], etc.) or are learned through a shrinkage function [78, 79]. Iterative solutions may be derived using splitting methods, such as half-quadratic splitting [80], reweighted least square [81], primal-dual[82], and alternating direction multiplier method [83]. Iterative methods have been applied in many coded aperture systems [84, 85, 86, 87, 88].
@@
Learning a direct mapping from $s$ to $\hat{b}$ using DNN has been extensively studied for many low-level computer vision tasks [105], including image deblurring, denoising, deraining, dehazing, in- and outpainting, as well as image restoration from low light conditions. Most vision models have an encoder-decoder framework, including UNet[106], residual network [107], and recurrent network [108]. The self-attention mechanism [109] seeks to capture long-range image dependencies. It serves as the building blocks of vision transformers [110, 111, 112]. Its quadratic computational complexities can be reduced to linear using window partitioning techniques [113, 114] or global approximation [115]. Self-attention has also been approximated by the spectral power density through the efficient fast Fourier transform (FFT) [116]. As an alternative, the spatial gating unit [117] was introduced with the MLP to achieve a performance comparable to that of transformers, where similar block partition and axis swap techniques were used to reduce its complexity [118]. Additionally, changing the image basis through linear or nonlinear transforms to better suit the orthogonally preconditioned optimizers (e.g., ADAM, ADAgrad, etc.) was found to accelerate the convergence and improve the performance of the DNNs. One notable finding involves the use of FFT features to reduce the spectral bias of neural networks and improve its learning of high-frequency information [119, 120, 121]. Another important achievement is the use of feature matching (or perceptual) loss [122] in the image intensity space, where features are extracted by pretrained VGG Net [123] or discriminators [124]. These feature representations and losses have been shown to increase the accuracy of image restoration [125, 126, 118, 127]. Furthermore, by breaking down a complicated task into subproblems and solving them progressively, multistage and multiscale frameworks allow supervision and feature fusion in multiple restoration stages and multiple image scales[128, 129] and encourage the recovery of image details. Embedding of kernel functions [130] and image coordinates [131, 132, 133] into neural networks has also been introduced, respectively, to utilize the knowledge of system PSF and positional information of an image.
@@ 4.1 Pre-Restoration with Conditional GAN
Here a residual UNet is designed for $G_{1}$, and the Markovian discriminator (PatchGAN) [146] is employed for $D_{1}$ in our SGAN-B/E models. The double convolutional layers of the original UNet are replaced with residual blocks [107]. The neural net blocks of this module are listed in Table II. For the SGAN-F model, we make use of the residual FFT (RFT) blocks [127] in $G_{1}$ and FFT convolutional (FTC) blocks in $D_{1}$ respectively. The structure of the net blocks and the values of the hyperparameters $\lambda_{ADV}$ and $\lambda_{GP}$ are provided in the Supplementary Information.
@@ 6.1 Baseline Methods and Evaluation Metrics
For the laser-dazzle protection, FFT representations and losses are found to be effective in restoring high-frequency image details. SSIM [173] and multiscale discriminator feature matching (MDF) losses [124] appear to have similar but weaker impacts on image structures restoration. L1 loss is shown to encourage in- and outpainting of missing areas, and VGG loss [122] improves perceptual qualities. Here we categorize the models as low- and high-performing based on the loss functions. The low-performing models, which trained with combinations of L1, SSIM, and MDF losses, involve MPRNet [129], Pix2Pix [146], ST-CGAN [172], Uformer [111], DeepWiener [92], WienerNet [91], and our SGAN-B. The high-performing models, which make use of FFT and VGG losses, include DeblurGAN [152], Stripformer [126], MIMO-UNet [125], MAXIM-2S [118], DeepRFT [127], and our SGAN-E/F. Neural networks exhibit a bias towards learning low frequency functions. The use of FFT features in DeepRFT [127] and our SGAN-F further improves their performances by suppressing the spectral bias.
@@ 6.6 Discussions
Future work is required to extend this approach to broadband illumination and a variety of laser wavelengths. Examination of other unfavorable imaging conditions, such as shift-variant distortions, motion blur, atmosphere, and adverse weathers, are also indispensable towards practical applications. Taking advantage of spectral representations in neural networks may be worth further investigation to develop fully blind image restoration without compromising the quality of the restored images. Aliasing issue presented in image downsampling may be further mitigated using learnable subpixel sampling techniques [179]. Provided limited computational resources, deep learning based super-resolution may allow restoration of high-resolution images from degraded inputs of low resolutions. Furthermore, a fully differentiable end-to-end imaging pipeline [180] that jointly optimizes the phase function and image restoration algorithm may further improve the performance of the antidazzle imaging system.